# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

### v1

@@ Appendix E Partially symmetric disorder
$\eta^{x}$ $\sim\mathcal{GP}(0,g^{2}C^{\phi}_{\star}),$ (146)
$\bm{H}=\begin{pmatrix}0&-1+g^{2}(S^{\phi}_{12})^{*}&0&0\\ -1+g^{2}S^{\phi}_{12}&-C^{\phi}_{12}&\sigma S^{\phi}_{1}C^{\phi}_{2}&\sigma S^{\phi}_{2}C^{\phi}_{1}\\ 0&\sigma(S^{\phi}_{1})^{*}C^{\phi}_{2}&0&\sigma-\sigma^{2}(S^{\phi}_{1})^{*}S^{\phi}_{2}\\ 0&\sigma(S^{\phi}_{2})^{*}C^{\phi}_{1}&\sigma-\sigma^{2}S^{\phi}_{1}(S^{\phi}_{2})^{*}&0\end{pmatrix}.$ (147)
Appendix F Inverse formula
In Sec. III.5.2 and Appendix H, we use the fact that
$\begin{pmatrix}0&a^{*}&0&-1\\ a&b&-1&0\\ 0&-1&0&c^{*}\\ -1&0&c&d\end{pmatrix}^{-1}=\begin{pmatrix}-\frac{b+|c|^{2}d}{|1-ac|^{2}}&-\frac{c}{1-ac}&-\frac{a^{*}d+bc}{|1-ac|^{2}}&-\frac{1}{1-ac}\\ -\frac{c^{*}}{1-(ac)^{*}}&0&-\frac{1}{1-(ac)^{*}}&0\\ -\frac{ad+bc^{*}}{|1-ac|^{2}}&-\frac{1}{1-ac}&-\frac{|a|^{2}d+b}{|1-ac|^{2}}&-\frac{a}{1-ac}\\ -\frac{1}{1-(ac)^{*}}&0&-\frac{a^{*}}{1-(ac)^{*}}&0\end{pmatrix}.$ (149)
Appendix G Two-site cavity calculation of the four-point function in the random-mode model
@@ G.1 Neuronal cavity
G.2 Latent-variable cavity
@@
[…]
@@
[…]
@@
$\gamma_{\alpha}(t)$ $=\sum_{i}R_{i\alpha}\phi_{i}(t),$ (167)
$G_{\alpha\beta}(t_{1},t_{2})$ $=\sqrt{N}\sum_{i,j}R_{i\alpha}L_{j\beta}S^{\phi}_{ij}(t_{1},t_{2}).$ (168)
@@
$C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega)=\sum_{i,j}R_{i\hat{0}}R_{j\hat{0}^{\prime}}C^{\phi}_{ij}(\omega),$ (169)
@@
$\Gamma_{G_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=\left\langle G_{\hat{0}\hat{0}^{\prime}}(\omega_{1})G_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\scriptsize\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (172)
$\Gamma_{G^{*}_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}^{\prime}\hat{0}}}(\bm{\omega})$ $=\left\langle G^{*}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})G_{\hat{0}^{\prime}\hat{0}}(\omega_{2})\right\rangle_{\scriptsize\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (173)
$\Gamma_{G^{*}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=\sqrt{N}\left\langle G^{*}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\scriptsize\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (174)
$\Gamma_{C^{\gamma}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=N\left\langle C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\scriptsize\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}}.$ (175)
@@
G.3 Combining the two two-site cavity pictures
@@
Using these relations and the definitions of the $\Gamma_{\cdots}(\bm{\omega})$ functions (Eqs. 159-162 and 172175), we obtain
@@
Appendix H Single-unit heterogeneity calculations
The full set of order parameters is
[…]
The action and single-site path integrals for this system are
$W^{z}_{D}[C^{\Phi},\hat{C}^{\Phi}]$ $=\int\mathcal{D}z\int\mathcal{D}\hat{z}\exp\left(i\hat{z}z-\frac{1}{2}z\hat{C}^{\Phi}z-\frac{1}{2}D^{2}\hat{z}C^{\Phi}\hat{z}\right).$ (195)
$\hat{C}^{\Phi}_{\star}(t_{1},t_{2})$ $=\hat{Q}_{\star}(t_{1},t_{2})=0,$ (196)
$C^{\Phi}_{\star}(t_{1},t_{2})$ $=\left\langle\left\langle\Phi_{\bm{\theta}}(t)\Phi_{\bm{\theta}}(t^{\prime})\right\rangle_{\star}\right\rangle_{\bm{\theta}},$ (197)
where $\left\langle\cdots\right\rangle_{\star}$ and $\left\langle\cdots\right\rangle_{\star D}$ are averages with respect to $W^{x}_{\bm{\theta}}[Q_{\star},0;0]$ and $W^{z}_{D}[C^{\Phi}_{\star},0]$, respectively (note that $W^{x}_{\bm{\theta}}[Q_{\star},0;0]$ does not depend on $\bm{\theta}$). Using $\left\langle z(t_{1})z(t_{2})\right\rangle_{\star D}=D^{2}C^{\Phi}_{\star}(t_{1},t_{2})$ and, consequently, $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\Phi}_{\star}(t_{1},t_{2})$, the self-consistent problem that determines $C^{\Phi}_{\star}(t_{1},t_{2})$ is given by Eq. 89 and the preceding description in Sec. VI.
The frequency-dependent Hessian at the saddle point is
$\bm{H}=\left(\begin{array}[]{cc|cc}0&\alpha r_{2}&0&-1\\ \alpha r_{2}&-\alpha r_{4}{C_{{12}}^{\Phi}}&-1&0\\ \hline\cr 0&-1&0&(\mathfrak{s}^{\Phi}_{12})^{*}\\ -1&0&\mathfrak{s}^{\Phi}_{12}&-\mathfrak{C}^{\Phi}_{12}\end{array}\right),$ (199)
where the new $\mathfrak{F}$raktur variables are defined in Eq. 90. We invert the Hessian using Eq. 149 and identify the upper-left element as Eq. 91. Combining the inverted Hessian with the relevant quantities for computing fluctuations of normalized variables,
$\frac{\delta^{2}S}{\delta\mathcal{J}^{2}}=-C^{\phi}_{12},\>\>\>\frac{\delta^{2}S}{\delta\bm{C}\delta\mathcal{J}}=\begin{pmatrix}0\\ 0\\ (S^{\phi}_{12})^{*}\\ -\mathfrak{C}^{\Phi\phi}_{12}\end{pmatrix},$ (200)
where
we obtain Eq. 92.
Appendix I Structured mode overlaps calculations
To generate matrices $\bm{L}$ and $\bm{R}$ with the correlation structure of Eq. 106 and Eq. 107, we express them in terms of independent Gaussian random matrices $\bm{X}^{1},\bm{X}^{2},$ and $\bm{Y}$, each with variance $1/N$:
${L}_{ia}$ $=\sqrt{{1-|\rho_{a}|}}{X}^{1}_{ia}+\sqrt{{|\rho_{a}|}}{Y}_{ia},$ (203)
[…]
$\centering Z=\int\mathcal{D}C^{\phi}\int\mathcal{D}\hat{C}^{\phi}\int\mathcal{D}S^{\phi}\@add@centering$ $\int\mathcal{D}\hat{S}^{\phi}\int\mathcal{D}Q\int\mathcal{D}\hat{Q}\int\mathcal{D}R\int\mathcal{D}\hat{R}\exp\left(-N\mathcal{S}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi},Q,\hat{Q},R,\hat{R}]\right),$ (209)
$W^{x}[Q,\hat{Q},R,\hat{R}]$ $=\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\left(i\hat{x}T[x]-\frac{1}{2}\hat{x}Q\hat{x}-\frac{1}{2}\phi\hat{Q}\phi-\frac{i}{2}\hat{x}R\phi-\frac{i}{2}\phi\hat{R}\hat{x}\right),$ (212)
$W^{z}_{D,\rho}[C^{\phi},\hat{C}^{\phi},S^{\phi},\hat{S}^{\phi}]$ $=\int\mathcal{D}z\int\mathcal{D}\hat{z}\exp\left(i\hat{z}z-\frac{1}{2}D^{2}\hat{z}C^{\phi}\hat{z}-\frac{1}{2}z\hat{C}^{\phi}z-\frac{i}{2}D\rho\hat{z}S^{\phi}z-\frac{i}{2}D\rho z\hat{S}^{\phi}\hat{z}\right).$ (213)
The saddle-point conditions yield:
$\hat{C}^{\phi}(t_{1},t_{2})$ $=\hat{Q}_{\star}(t_{1},t_{2})=0,$ (222)
$S^{\phi}_{\star}(t_{1},t_{2})$ $=\hat{S}^{\phi}_{\star}(t_{1},t_{2})=\left\langle\frac{\delta\phi(t_{1})}{\delta I(t_{2})}\right\rangle_{\star},$ (224)
$R_{\star}(t_{1},t_{2})$ $=\hat{R}_{\star}(t_{1},t_{2})=\alpha\left\langle D\rho\left\langle\frac{\delta z(t_{1})}{\delta I(t_{2})}\right\rangle_{\star D,\rho}\right\rangle_{D,\rho},$ (226)
$T[x](t)$ $=\eta^{x}(t)+[R_{\star}\ast\phi](t)$ (227)
$z(t)$ $=\eta^{z}(t)+D\rho[S^{\phi}_{\star}\ast z](t),$ (228)
where the Gaussian fields have statistics
$\eta^{x}$ $\sim\mathcal{GP}(0,Q_{\star}),$ (229)
$\eta^{z}$ $\sim\mathcal{GP}(0,D^{2}C^{\phi}_{\star}).$ (230)
This gives rise to the single-site problem described in Sec. VII.
The frequency-dependent Hessian at the saddle point is
$\bm{H}=\left(\begin{array}[]{cccc|cccc}0&u^{*}&0&0&0&-1&0&0\\ u&-vC^{\phi}_{1}C^{\phi}_{2}&x_{12}C_{2}&x_{21}C^{\phi}_{1}&-1&0&0&0\\ 0&x_{12}^{*}C^{\phi}_{2}&0&-w&0&0&0&1\\ 0&x_{21}^{*}C^{\phi}_{1}&-w^{*}&0&0&0&1&0\\ \hline\cr 0&-1&0&0&0&(S^{\phi}_{1}S^{\phi}_{2})^{*}&0&0\\ -1&0&0&0&S^{\phi}_{1}S^{\phi}_{2}&-C^{\phi}_{1}C^{\phi}_{2}&S^{\phi}_{1}C^{\phi}_{2}&S^{\phi}_{2}C^{\phi}_{1}\\ 0&0&0&1&0&(S^{\phi}_{1})^{*}C^{\phi}_{2}&0&-(S^{\phi}_{1})^{*}S^{\phi}_{2}\\ 0&0&1&0&0&(S^{\phi}_{2})^{*}C^{\phi}_{1}&-S^{\phi}_{1}(S^{\phi}_{2})^{*}&0\end{array}\right)$ (231)
where
$\sigma_{k}$ $=\frac{1}{1-D\rho S^{\phi}_{k}}\>\>\text{for}\>\>k=1,2.$ (236)
Appendix J Numerical details
J.1 Numerics for theory
To validate the theory, we used networks with dynamics defined by $T[x](t)=(1+\partial_{t})x(t)$ (Eq. 2) and with nonlinearity $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$. For all connectivity models, except those with structured $\bm{L}$-$\bm{R}$ overlaps (VII), the single-site two-point functions $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$ depend only on $g_{\text{eff}}$ and can be numerically calculated as in i.i.d. networks (see, e.g., [10, 77]). In summary:
@@ Appendix I Structured mode overlaps calculations
    Squaring the single-site picture gives a second-order ODE $\partial^{2}_{\tau}C^{x}_{\star}(\tau)=C^{x}_{\star}(\tau)-g_{\text{eff}}^{2}C^{\phi}_{\star}(\tau)$. Since the rhs depends only on $C^{x}_{\star}(\tau)$ and $C^{x}_{\star}(0)$, we can consider it to be the negative derivative $-\partial V/\partial C^{x}_{\star}(\tau)$ of a potential $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$ with an explicit dependence on the initial condition $C^{x}_{\star}(0)$.
@@
    Restricting to solutions with $C^{x}_{\star}(\tau)\rightarrow 0$ as $\tau\rightarrow\infty$ and $\partial_{\tau}C^{x}_{\star}(\tau)|_{\tau=0}=0$, enforcing conservation of energy gives $V(C^{x}_{\star}(0),C^{x}_{\star}(0))=V(0,C^{x}_{\star}(0))$, which can be solved for $C^{x}_{\star}(0)$ by numerical root finding.
@@
    Finally, Euler integration of the $[C^{x}_{\star}(\tau),\partial_{\tau}C^{x}_{\star}(\tau)]$ dynamics with these initial conditions gives its full time course, along with $C^{\phi}_{\star}(\tau)$ via Eq. 237.
We integrated the Newtonian ODE from $\tau=0$ to $\tau_{\text{max}}=200$ with step size $d\tau=0.025$. When then use an FFT to otain the frequency-space representation. The frequency-space representation of the response function $S^{\phi}_{\star}(\omega)$ was computed directly as
@@
    Each of $N_{\text{sample}}$ single-site processes $x_{s}(t)$ is Euler integrated according to Eq. 110 with the input $\eta_{s}^{x}(t)$ and self-feedback by the convolution between $R_{0}(\tau)$ and $\phi(x_{s}(t))$.

### v2

@@ I.2 Four-point functions
$R_{\star}(t_{1},t_{2})$ $=\hat{R}_{\star}(t_{2},t_{1})=\alpha\left\langle D\rho\left\langle\frac{\delta z(t_{1})}{\delta I(t_{2})}\right\rangle_{\star D,\rho}\right\rangle_{D,\rho},$ (188)
$T[x](t)$ $=\eta^{x}(t)+[R_{\star}\ast\phi](t)$ (189)
$z(t)$ $=\eta^{z}(t)+D\rho[S^{\phi}_{\star}\ast z](t),$ (190)
where the Gaussian fields have statistics
$\eta^{x}$ $\sim\mathcal{GP}(0,Q_{\star}),$ (191)
$\eta^{z}$ $\sim\mathcal{GP}(0,D^{2}C^{\phi}_{\star}).$ (192)
This gives rise to the single-site problem stated in Sec. VII.
K.2 Four-point functions
The frequency-dependent Hessian at the saddle point is
$\bm{H}=\left(\begin{array}[]{cccc|cccc}0&u&0&0&0&-1&0&0\\ u^{*}&-vC^{\phi}_{1}C^{\phi}_{2}&x^{*}_{12}C^{\phi}_{2}&x^{*}_{21}C^{\phi}_{1}&-1&0&0&0\\ 0&x_{12}C^{\phi}_{2}&0&-w^{*}&0&0&0&1\\ 0&x_{21}C^{\phi}_{1}&-w&0&0&0&1&0\\ \hline\cr 0&-1&0&0&0&S^{\phi}_{1}S^{\phi}_{2}&0&0\\ -1&0&0&0&(S^{\phi}_{1}S^{\phi}_{2})^{*}&-C^{\phi}_{1}C^{\phi}_{2}&(S^{\phi}_{1})^{*}C^{\phi}_{2}&(S^{\phi}_{2})^{*}C^{\phi}_{1}\\ 0&0&0&1&0&S^{\phi}_{1}C^{\phi}_{2}&0&-S^{\phi}_{1}(S^{\phi}_{2})^{*}\\ 0&0&1&0&0&S^{\phi}_{2}C^{\phi}_{1}&-(S^{\phi}_{1})^{*}S^{\phi}_{2}&0\end{array}\right)$ (193)
where
$\sigma_{k}$ $=\frac{1}{1-D\rho S^{\phi}_{k}}\>\>\text{for}\>\>k=1,2.$ (198)
K.3 Sigmoidal parameterization of overlaps
To model structured left–right mode overlaps, we used the following sigmoidal form:
$\rho_{a}=\gamma_{\rho}\left(1-\frac{2}{1+\exp(-\beta_{\rho}(a/N-u_{\text{flip}}))}\right),$ (199)
where $\rho_{a}=0$ at $a/N=u_{\text{flip}}$. The parameter $\gamma_{\rho}$ sets the overall magnitude and sign of the overlaps, while $\beta_{\rho}$ controls the steepness of the transition. We fixed $\beta_{\rho}=16$, varied $\gamma_{\rho}$, and set $u_{\text{flip}}$ such that $\left\langle D\rho\right\rangle_{D,\rho}=0$, thereby eliminating strong self-couplings.
Appendix L Two-site cavity calculation of four-point functions for the random-mode model
While the path-integral approach provides a systematic framework for analyzing neural-network dynamics, the cavity method offers an intuitive, complementary perspective. This section provides an overview of the cavity analysis for the random-mode model, with the full mathematical details available in Appendix L. While these approaches yield equivalent results, they differ in their methodology: the path-integral approach focuses on fluctuations around the saddle point in a field theory, whereas the cavity approach examines perturbations induced by a subset of held-out variables. Here we compute the four-point function using a two-site cavity approach, noting that the two-point function (which, if desired, could be computed using a single-site cavity method) is the same as that of an i.i.d. network with appropriate $g_{\text{eff}}$.
Leveraging Eq. 125, we reformulate the network as a bipartite system of neurons and latent variables and extend the cavity calculation of Clark et al. [30] to both groups. This leads to separate, mutually referential cavity pictures for each group, mirroring the two single-site path integrals (Sec. IV). As in the path-integral formalism, self-consistent expressions in one picture are defined using averages from the other. Unlike the path-integral approach, which derives $\Psi^{\phi}(\bm{\tau})$ using its time-by-time definition (Eq. 25), the cavity method derives this function using its neuron-by-neuron definition (Eq. 24).
The cavity method we employ, schematized in Fig. 11, has two key features that distinguish it from simpler cavity calculations. First, it uses a two-site structure: we consider the simultaneous removal of two neurons, or two latent variables, from the network, rather than just one. This allows us to study cross-correlations between neurons. This approach was previously used in Clark et al. [30]. Second, it incorporates a bipartite structure: we treat both neurons and latent variables as dynamic objects, creating separate but mutually referential cavity pictures for each group. Such bipartite structure, without the two-site structure, has been used in cavity calculations in the Hopfield model [96, 97]. Note that these two features are distinct: the two-site aspect refers to the number of neurons removed in each cavity, while the bipartite aspect refers to the types of variables considered (neurons and latent variables).
@@
L.2 Latent-variable cavity
@@
$\delta\phi_{i}(t)=\int^{t}dt^{\prime}\sum_{j=1}^{N}S^{\phi}_{ij}(t_{1},t_{2})\sum_{\alpha}L_{j\alpha}z_{\alpha}(t^{\prime}),$ (215)
@@
$z_{\alpha}(t)=D_{\alpha}\left[\gamma_{\alpha}(t)+\frac{1}{\sqrt{N}}\sum_{\beta}\int^{t}dt^{\prime}G_{\alpha\beta}(t_{1},t_{2})z_{\beta}(t^{\prime})\right],$ (216)
@@
$\gamma_{\alpha}(t)$ $=\sum_{i=1}^{N}R_{i\alpha}\phi_{i}(t),$ (217)
$G_{\alpha\beta}(t_{1},t_{2})$ $=\sqrt{N}\sum_{i,j=1}^{N}R_{i\alpha}L_{j\beta}S^{\phi}_{ij}(t_{1},t_{2}).$ (218)
@@
$C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega)=\sum_{i,j=1}^{N}R_{i\hat{0}}R_{j\hat{0}^{\prime}}C^{\phi}_{ij}(\omega),$ (219)
@@
$\Gamma_{G_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=\left\langle G_{\hat{0}\hat{0}^{\prime}}(\omega_{1})G_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (222)
$\Gamma_{G^{*}_{\hat{0}\hat{0}^{\prime}}G_{\hat{0}^{\prime}\hat{0}}}(\bm{\omega})$ $=\left\langle G^{*}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})G_{\hat{0}^{\prime}\hat{0}}(\omega_{2})\right\rangle_{\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (223)
$\Gamma_{G^{*}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=\sqrt{N}\left\langle G^{*}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}},$ (224)
$\Gamma_{C^{\gamma}_{\hat{0}\hat{0}^{\prime}}C^{\gamma}_{\hat{0}\hat{0}^{\prime}}}(\bm{\omega})$ $=N\left\langle C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{1})C^{\gamma}_{\hat{0}\hat{0}^{\prime}}(\omega_{2})\right\rangle_{\begin{subarray}{l}\bm{L},\bm{R},\\ \bm{D}\setminus\{D_{\hat{0}},D_{\hat{0}^{\prime}}\}\end{subarray}}.$ (225)
@@
L.3 Combining the two two-site cavity pictures
@@
Using these relations and the definitions of the $\Gamma_{\cdots}(\bm{\omega})$ functions (Eqs. 209-212 and 222225), we obtain
@@
Appendix M Numerical details
M.1 Numerics for theory
To validate the theory, we used networks with dynamics defined by $T[x](t)=(1+\partial_{t})x(t)$ and with nonlinearity $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$. For all connectivity models, except those with structured $\bm{L}$-$\bm{R}$ overlaps (Sec. VII), the single-site two-point functions $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$ depend only on $g_{\text{eff}}$ and can be numerically calculated as in i.i.d. networks (see, e.g., [10, 98]). In summary:
@@
    Squaring the single-site picture gives a second-order ODE $\partial^{2}_{\tau}C^{x}_{\star}(\tau)=C^{x}_{\star}(\tau)-g_{\text{eff}}^{2}C^{\phi}_{\star}(\tau)$. Since the rhs depends only on $C^{x}_{\star}(\tau)$ and $C^{x}_{\star}(0)$, we can consider it to be the negative derivative $\partial V/\partial C^{x}_{\star}(\tau)$ of a potential $V(C^{x}_{\star}(\tau),C^{x}_{\star}(0))$ with an explicit dependence on the initial condition $C^{x}_{\star}(0)$.
@@
    Restricting to solutions with $C^{x}_{\star}(\tau)\to 0$ as $\tau\to\infty$ and $\partial_{\tau}C^{x}_{\star}(\tau)|_{\tau=0}=0$, enforcing conservation of energy gives $V(C^{x}_{\star}(0),C^{x}_{\star}(0))=V(0,C^{x}_{\star}(0))$, which can be solved for $C^{x}_{\star}(0)$ by numerical root finding.
@@
    Finally, Euler integration of the $C^{x}_{\star}(\tau),\partial_{\tau}C^{x}_{\star}(\tau)]$ dynamics with these initial conditions gives its full time course, along with $C^{\phi}_{\star}(\tau)$ via Eq. 241.
We integrated the Newtonian ODE from $\tau=0$ to $\tau_{\text{max}}=200$ with step size $d\tau=0.025$. When then use an FFT to obtain the frequency-space representation. The frequency-space representation of the response function $S^{\phi}_{\star}(\omega)$ was computed directly as
@@
    Each of $N_{\text{sample}}$ single-site processes $x_{s}(t)$ is Euler integrated according to Eq. 47a with the input $\eta_{s}^{x}(t)$ and self-feedback by the convolution between $R_{0}(\tau)$ and $\phi(x_{s}(t))$.